OpenStudy (anonymous):

A store had 250 bottles of water. Each week, 40% of the bottles were sold and 48 new bottles arrived in shipments. Which recursive function best represents the number of bottles of water in the store?

OpenStudy (anonymous):

@Kira_Yamato

OpenStudy (anonymous):

f(n) = f(n - 1) • 0.4 + 48, n > 0 f(n) = 250 - f(n - 1) • 0.4 + 48, n > 0 f(n) = f(n - 1) • 0.6 + 48, n > 0 f(n) = 250 - f(n - 1) • 0.6 + 48, n > 0

OpenStudy (anonymous):

I think it's A or B.

OpenStudy (kira_yamato):

A store had 250 bottles of water. => f(0) = 250 Each week, 40% of the bottles were sold and 48 new bottles arrived in shipments. => \[f(n) = 0.4f(n-1) + 48, n \in \mathbb{Z^+}\]

OpenStudy (anonymous):

So the answer is A?